# EUROPEAN ORGANIZATION FOR NUCLEAR RESEARCH (CERN)

### v1

@@
[Uncaptioned image] CERN-EP-2025-171 LHCb-PAPER-2025-026 August 18, 2025
@@
An amplitude analysis of the ${{{B}^{+}}}\rightarrow({{\pi}^{+}}{{\pi}^{-}})({{K}^{0}_{\mathrm{S}}}{{\pi}^{+}})$ decay is performed in the mass regions $0.30<m_{{{\pi}^{+}}{{\pi}^{-}}}<1.10\text{\,Ge\kern-1.00006ptV\!/}c^{2}$ and $0.75<m_{{{K}^{0}_{\mathrm{S}}}{{\pi}^{+}}}<1.20\text{\,Ge\kern-1.00006ptV\!/}c^{2}$, using $pp$ collision data recorded with the LHCb detector corresponding to an integrated luminosity of $9\text{\,fb}^{-1}$. The polarization fractions and CP{C\!P} asymmetries for ${{{B}^{+}}}\rightarrow\rho(770)^{0}{{K}^{*}}(892)^{+}$ decays are measured. Violation of the CP{C\!P} symmetry in the decay ${{{B}^{+}}}\rightarrow\rho(770)^{0}{{K}^{*}}(892)^{+}$ is observed for the first time, with a significance exceeding nine standard deviations. The CP{C\!P} asymmetry is measured to be $\mathcal{A}_{{C\!P}}={0.507\pm 0.062\text{\,(stat)}\pm 0.017\text{\,(syst)}}$ and the CP{C\!P}-averaged longitudinal polarization fraction of $f_{L}={0.720\pm 0.028\text{\,(stat)}\pm 0.009\text{\,(syst)}}$. The measurements help to shed light on the polarization puzzle of $B$ mesons decaying to two vector mesons.
Submitted to Phys. Rev. Lett.
© 2025 CERN for the benefit of the LHCb collaboration. CC BY 4.0 licence.
@@
The specific class of ${B}\rightarrow VV$ decays with ${B}\rightarrow\rho{{K}^{*}}$ has been widely studied, where, throughout this Letter, $\rho$ and ${K}^{*}$ represent the $\rho(770)$ and ${{K}^{*}}(892)$ mesons, respectively, unless otherwise stated. The longitudinal fractions are measured to be substantially different depending on the final state, with $f_{L}(\mbox{${{{B}^{+}}}\!\rightarrow\rho^{0}{{K}^{*}}^{+}$})=$ $0.78\pm 0.12$ [37], $f_{L}({{{B}^{+}}}\rightarrow{{\rho}^{+}}{{K}^{*0}})=0.48\pm 0.08$ [38, 39, 5], $f_{L}({{B}^{0}}\rightarrow{{\rho}^{0}}{{K}^{*0}})=$ $0.173\pm 0.026$ [38, 40, 15], and $f_{L}({{B}^{0}}\rightarrow{{\rho}^{-}}{{K}^{*+}})=0.38\pm 0.13$ [40], where charge conjugation is implied throughout this Letter unless otherwise specifically stated. These observed discrepancies reflect the varying interference patterns between tree-level and loop processes [41], which would also be revealed by measuring the CPC\!P asymmetries. Among these decays, CPC\!P violation has only been established in the longitudinal component of the ${{B}^{0}}\rightarrow{{\rho}^{0}}{{K}^{*0}}$ decay [15]. Detailed and precise measurements of ${B}\rightarrow\rho{{K}^{*}}$ decays, compared with various predictions [2, 20, 21, 22, 24, 25, 41, 42, 43, 44, 45], permit more precise extractions of nonperturbative parameters through global fits [45] and help to shed light on the polarization puzzle and the dynamics of CPC\!P violation in these decays. They also comprise fundamental input for theoretically clean, model-independent isospin sum rules that are highly sensitive to physics beyond the SM [46].
@@
$\propto$ $\Phi\cdot\left|\sum_{i=1}^{N}A_{i}\cdot g_{i}\bigl{(}\cos\theta_{{{\pi}^{+}}{{\pi}^{-}}},\cos\theta_{{{K}^{0}_{\mathrm{S}}}{{\pi}^{+}}},\phi\bigr{)}\cdot M_{i}\bigl{(}m_{{{\pi}^{+}}{{\pi}^{-}}},m_{{{K}^{0}_{\mathrm{S}}}{{\pi}^{+}}}\bigr{)}\right|^{2},$
@@
where $\Phi$ is the four-body phase-space density and $A_{i}$ is the complex coupling for each amplitude. For the ${B}^{-}$}} decay, the sign of the $\phi$ angle is flipped and an independent set of complex couplings, $\kern 1.79993pt\overline{\kern-1.79993ptA}_{i}$, is measured. The $g_{i}$ functions are characterized by spherical harmonics describing the angular distribution. The twenty-two amplitudes included in the baseline fit are detailed in Table 3 of the Supplemental Material [66]. The $M_{i}$ functions describe the mass distributions of the ${{\pi}^{+}}{{\pi}^{-}}$ and ${{K}^{0}_{\mathrm{S}}}{{\pi}^{+}}$ systems. They are briefly described below and detailed in Table 4 of the Supplemental Material [66].
@@
In the ${{\pi}^{+}}{{\pi}^{-}}$ mass spectrum, besides the $\rho(770)^{0}$ meson, a few other resonances are identified, including the $\omega(782)$, $f_{0}(500)$, $f_{0}(980)$ and $f_{0}(1370)$ mesons. The mass distributions of the vector resonances, $\rho(770)^{0}$ and $\omega(782)$, are modeled by a Gounaris–Sakurai parameterization [67] and a relativistic Breit–Wigner function, respectively. Those for the scalar resonances, $f_{0}(500)$ and $f_{0}(1370)$, are described by relativistic Breit–Wigner functions, while that for the $f_{0}(980)$ meson is modeled by a Flatté parameterization [68, 69, 70]. The ${{K}^{0}_{\mathrm{S}}}{{\pi}^{+}}$ mass distribution is dominated by the ${{K}^{*}}(892)^{+}$ resonance and a scalar contribution, referred to as the $({{K}^{0}_{\mathrm{S}}}{{\pi}^{+}})_{\text{S}}$ component. The ${{K}^{*}}(892)^{+}$ resonance is described by a relativistic Breit–Wigner. The $({{K}^{0}_{\mathrm{S}}}{{\pi}^{+}})_{\text{S}}$ distribution is parameterized by the LASS [71] model, which includes the contributions of the $K^{*}_{0}(1430)^{+}$ resonance and a broad low-mass component. A small fraction of ${{{B}^{+}}}\rightarrow a_{1}(1260)^{+}{{K}^{0}_{\mathrm{S}}}$ and ${{{B}^{+}}}\rightarrow a_{1}(1640)^{+}{{K}^{0}_{\mathrm{S}}}$ decays is observed in the sample, with $a_{1}^{+}\rightarrow{{\rho}^{0}}(\rightarrow{{\pi}^{+}}{{\pi}^{-}}){{\pi}^{+}}$, $a_{1}^{+}\rightarrow f_{0}(500)(\rightarrow{{\pi}^{+}}{{\pi}^{-}}){{\pi}^{+}}$, and $a_{1}^{+}\rightarrow f_{0}(980)(\rightarrow{{\pi}^{+}}{{\pi}^{-}}){{\pi}^{+}}$. The $a_{1}$ mass distributions are described by relativistic Breit–Wigner functions, with the angular variables redefined accordingly. Parameters describing the mass distributions are fixed in the baseline fit according to the results of previous measurements [38, 15].
@@
Two main categories of systematic uncertainty contribute to the result: those originating from experimental assumptions and those related to the amplitude model. The experimental category includes effects due to the limited precision of the background fractions, the efficiency map and background modeling. The amplitude model uncertainties arise from limited knowledge of the fixed parameters of the resonance mass models, and possible spin-2 components neglected in the baseline fit. Systematic uncertainties due to experimental effects are found to be subdominant. The breakdown of the systematic uncertainties of observables is detailed in Table End Matter of the End Matter. Results of the fitted parameters with their uncertainties are also summarized in Table 2 of the End Matter.
@@
The systematic uncertainty due to fixed parameters in the amplitude model is estimated using repeated fits to the data. Each time the masses, widths, and hadron radii of intermediate resonances are varied within their uncertainties and an amplitude fit is performed under this model. The standard deviation of the distribution of the fit results among the repeated fits is taken as the corresponding systematic uncertainty. For the $B\rightarrow VV$ decay, the relative orbital angular momentum of the two vector mesons, $L$, could take the values $L=0$ or $L=2$ for the $A_{0}$ and $A_{\parallel}$ amplitudes, and $L=1$ for the $A_{\perp}$ amplitude. Various $L$ configurations are studied and the results are compared with that of the baseline fit with $L=0$, and the largest difference is taken as the systematic uncertainty. Furthermore, possible contributions of spin-2 resonances to the decay are assessed by including the $f_{2}(1270)\rightarrow{{\pi}^{+}}{{\pi}^{-}}$ or $K_{2}^{*}(1430)^{+}\rightarrow{{K}^{0}_{\mathrm{S}}}{{\pi}^{+}}$ components in the fit. The systematic uncertainties are estimated using the mean value of deviations, between the fits using extended model and baseline model, over an ensemble of pseudoexperiments.
@@
$f_{L}^{+}\equiv\frac{|A_{0}|^{2}}{\sum_{\lambda}|A_{\lambda}|^{2}}={0.491\pm 0.083\pm 0.025},\quad f_{L}^{-}\equiv\frac{|\kern 1.79993pt\overline{\kern-1.79993ptA}_{0}|^{2}}{\sum_{\lambda}|\kern 1.79993pt\overline{\kern-1.79993ptA}_{\lambda}|^{2}}={0.794\pm 0.025\pm 0.007}.$
@@
$\mathcal{A}_{{C\!P}}\equiv\frac{\sum_{\lambda}(|\kern 1.79993pt\overline{\kern-1.79993ptA}_{\lambda}|^{2}-|A_{\lambda}|^{2})}{\sum_{\lambda}(|\kern 1.79993pt\overline{\kern-1.79993ptA}_{\lambda}|^{2}+|A_{\lambda}|^{2})}={0.507\pm 0.062\pm 0.017}.$
@@
$\mathcal{A}_{{C\!P}}(A_{0})$ $\equiv\frac{|\kern 1.79993pt\overline{\kern-1.79993ptA}_{0}|^{2}-|A_{0}|^{2}}{|\kern 1.79993pt\overline{\kern-1.79993ptA}_{0}|^{2}+|A_{0}|^{2}}$ $={0.664\pm 0.083\pm 0.025},$
@@
$\mathcal{A}_{{C\!P}}(A_{\parallel})$ $={-0.063\pm 0.137\pm 0.024},\phantom$$\mathrm${rad}}}
@@
The $\it{true}$ TPA $\mathcal{A}_{\it{true}}^{(1)}$ deviates from zero with a significance of approximately four standard deviations, indicating evidence of CPC\!P violation in interference. The $\it{fake}$ TPA $\mathcal{A}_{\it{fake}}^{(1)}$ is nonzero with a significance of six standard deviations, marking parity violation in interference [75].
In summary, an amplitude analysis of the decay ${{{B}^{+}}}\rightarrow({{\pi}^{+}}{{\pi}^{-}})({{K}^{0}_{\mathrm{S}}}{{\pi}^{+}})$ is performed within the mass ranges $0.30<m_{{{\pi}^{+}}{{\pi}^{-}}}<1.10\text{\,Ge\kern-1.00006ptV\!/}c^{2}$ and $0.75<m_{{{K}^{0}_{\mathrm{S}}}{{\pi}^{+}}}<1.20\text{\,Ge\kern-1.00006ptV\!/}c^{2}$. The longitudinal polarization and CPC\!P asymmetries are studied for the ${{{B}^{+}}}\rightarrow\rho(770)^{0}{{K}^{*}}(892)^{+}$ decay. The CPC\!P-averaged longitudinal fraction is measured to be $f_{L}={0.720\pm 0.028\text{\,(stat)}\pm 0.009\text{\,(syst)}}$. The direct CPC\!P asymmetry is determined to be $\mathcal{A}_{{C\!P}}={0.507\pm 0.062\text{\,(stat)}\pm 0.017\text{\,(syst)}}$, and the significance of CPC\!P violation is measured to exceed nine standard deviations using a likelihood-ratio test. This is the first observation of CPC\!P violation in the ${{{B}^{+}}}\rightarrow\rho(770)^{0}{{K}^{*}}(892)^{+}$ decay. These measurements are consistent with those previously reported by the BaBar collaboration [37], with much better precision. The CPC\!P asymmetries are also derived for the magnitudes and phases of the three subamplitudes, suggesting that the longitudinal component drives the total CPC\!P violation. These results help to constrain theoretical models used to calculate $B$-hadron decays, shed light on the dynamics of heavy-flavor decays and contribute to the understanding of the polarization puzzle of $B$-meson decays.

### v2

@@
[Uncaptioned image] CERN-EP-2025-171 LHCb-PAPER-2025-026 January 26, 2026
@@
An amplitude analysis of the ${{{B}^{+}}}\rightarrow({{\pi}^{+}}{{\pi}^{-}})({{K}^{0}_{\mathrm{S}}}{{\pi}^{+}})$ decay is performed in the mass regions $0.30<m_{{{\pi}^{+}}{{\pi}^{-}}}<1.10\text{\,Ge\kern-1.00006ptV\!/}c^{2}$ and $0.75<m_{{{K}^{0}_{\mathrm{S}}}{{\pi}^{+}}}<1.20\text{\,Ge\kern-1.00006ptV\!/}c^{2}$, using $pp$ collision data recorded with the LHCb detector corresponding to an integrated luminosity of $9\text{\,fb}^{-1}$. The polarization fractions and CP{C\!P} asymmetries for ${{{B}^{+}}}\rightarrow\rho(770)^{0}{{K}^{*}}(892)^{+}$ decays are measured. Violation of the CP{C\!P} symmetry in the decay ${{{B}^{+}}}\rightarrow\rho(770)^{0}{{K}^{*}}(892)^{+}$ is observed for the first time, with a significance exceeding nine standard deviations. The CP{C\!P} asymmetry is measured to be $\mathcal{A}_{{C\!P}}={0.507\pm 0.062\text{\,(stat)}\pm 0.024\text{\,(syst)}}$ and the CP{C\!P}-averaged longitudinal polarization fraction of $f_{L}={0.720\pm 0.028\text{\,(stat)}\pm 0.009\text{\,(syst)}}$. The measurements help to shed light on the polarization puzzle of $B$ mesons decaying to two vector mesons.
Published in Phys. Rev. Lett.
© 2026 CERN for the benefit of the LHCb collaboration. CC BY 4.0 licence.
@@
The specific class of ${B}\rightarrow VV$ decays with ${B}\rightarrow\rho{{K}^{*}}$ has been widely studied, where, throughout this Letter, $\rho$ and ${K}^{*}$ represent the $\rho(770)$ and ${{K}^{*}}(892)$ mesons, respectively, unless otherwise stated. The longitudinal fractions are measured to be substantially different depending on the final state, with $f_{L}(\mbox{${{{B}^{+}}}\!\rightarrow\rho^{0}{{K}^{*}}^{+}$})=$ $0.78\pm 0.12$ [37], $f_{L}({{{B}^{+}}}\rightarrow{{\rho}^{+}}{{K}^{*0}})=0.48\pm 0.08$ [38, 39, 5], $f_{L}({{B}^{0}}\rightarrow{{\rho}^{0}}{{K}^{*0}})=$ $0.173\pm 0.026$ [38, 40, 15], and $f_{L}({{B}^{0}}\rightarrow{{\rho}^{-}}{{K}^{*+}})=0.38\pm 0.13$ [40], where charge conjugation is implied throughout this Letter unless otherwise specifically stated. These observed discrepancies reflect the varying interference patterns between tree-level and loop processes [41], which would also be revealed by measuring the CPC\!P asymmetries. Among these decays, CPC\!P violation has only been established in the longitudinal component of the ${{B}^{0}}\rightarrow{{\rho}^{0}}{{K}^{*0}}$ decay [15]. Detailed and precise measurements of ${B}\rightarrow\rho{{K}^{*}}$ decays, compared with various predictions [2, 20, 21, 22, 24, 25, 41, 42, 43, 44, 45, 46], permit more precise extractions of nonperturbative parameters through global fits [46] and help to shed light on the polarization puzzle and the dynamics of CPC\!P violation in these decays. They also comprise fundamental input for theoretically clean, model-independent isospin sum rules that are highly sensitive to physics beyond the SM [47].
@@
$\propto$ $\Phi\times\left|\sum_{i=1}^{N}A_{i}\times g_{i}\bigl(\cos\theta_{{{\pi}^{+}}{{\pi}^{-}}},\cos\theta_{{{K}^{0}_{\mathrm{S}}}{{\pi}^{+}}},\phi\bigr)\times M_{i}\bigl(m_{{{\pi}^{+}}{{\pi}^{-}}},m_{{{K}^{0}_{\mathrm{S}}}{{\pi}^{+}}}\bigr)\right|^{2},$
@@
where $\Phi$ is the four-body phase-space density and $A_{i}$ is the complex coupling for each amplitude. By construction, the interference terms between components with different spin–parity quantum numbers cancel when integrating over the phase space. For the ${B}^{-}$}} decay, the sign of the $\phi$ angle is flipped and an independent set of complex couplings, $\kern 1.79993pt\overline{\kern-1.79993ptA}_{i}$, is measured. The $g_{i}$ functions are characterized by spherical harmonics describing the angular distribution. The twenty-two amplitudes included in the baseline fit are detailed in the Supplemental Material [67]. The $M_{i}$ functions describe the mass distributions of the ${{\pi}^{+}}{{\pi}^{-}}$ and ${{K}^{0}_{\mathrm{S}}}{{\pi}^{+}}$ systems. They are briefly described below and detailed in the Supplemental Material [67].
@@
In the ${{\pi}^{+}}{{\pi}^{-}}$ mass spectrum, besides the $\rho(770)^{0}$ meson, a few other resonances are identified, including the $\omega(782)$, $f_{0}(500)$, $f_{0}(980)$ and $f_{0}(1370)$ mesons. The mass distributions of the vector resonances, $\rho(770)^{0}$ and $\omega(782)$, are modeled by a Gounaris–Sakurai parameterization [68] and a relativistic Breit–Wigner function, respectively. Those for the scalar resonances, $f_{0}(500)$ and $f_{0}(1370)$, are described by relativistic Breit–Wigner functions, while that for the $f_{0}(980)$ meson is modeled by a Flatté parameterization [69, 70, 71]. Several fits were performed with and without the scalar $f_{0}$ states; all states included in the baseline fit have an individual significance above six standard deviations. Alternative descriptions of the $\pi\pi$ S-wave using the K-matrix formalism are considered as a source of systematic uncertainty, which is detailed below. The ${{K}^{0}_{\mathrm{S}}}{{\pi}^{+}}$ mass distribution is dominated by the ${{K}^{*}}(892)^{+}$ resonance and a scalar contribution, referred to as the $({{K}^{0}_{\mathrm{S}}}{{\pi}^{+}})_{\text{S}}$ component. The ${{K}^{*}}(892)^{+}$ resonance is described by a relativistic Breit–Wigner. The $({{K}^{0}_{\mathrm{S}}}{{\pi}^{+}})_{\text{S}}$ distribution is parameterized by the LASS [72] model, which includes the contributions of the $K^{*}_{0}(1430)^{+}$ resonance and a broad low-mass component. Possible $({{K}^{0}_{\mathrm{S}}}{{\pi}^{+}})_{\text{P}}$ contributions from the $K_{1}(1270)$, $K_{1}(1400)$ and $K_{1}^{*}(1410)$ states were considered, but no significant effects are expected in the studied mass range. A small fraction of ${{{B}^{+}}}\rightarrow a_{1}(1260)^{+}{{K}^{0}_{\mathrm{S}}}$ and ${{{B}^{+}}}\rightarrow a_{1}(1640)^{+}{{K}^{0}_{\mathrm{S}}}$ decays is observed in the sample, with $a_{1}^{+}\rightarrow{{\rho}^{0}}(\rightarrow{{\pi}^{+}}{{\pi}^{-}}){{\pi}^{+}}$, $a_{1}^{+}\rightarrow f_{0}(500)(\rightarrow{{\pi}^{+}}{{\pi}^{-}}){{\pi}^{+}}$, and $a_{1}^{+}\rightarrow f_{0}(980)(\rightarrow{{\pi}^{+}}{{\pi}^{-}}){{\pi}^{+}}$. The $a_{1}$ mass distributions are described by relativistic Breit–Wigner functions, with the angular variables redefined accordingly. Parameters describing the mass distributions are fixed in the baseline fit according to the results of previous measurements [38, 15].
@@
Two main categories of systematic uncertainty contribute to the result: those originating from experimental assumptions and those related to the amplitude model. The experimental category includes effects due to the limited precision of the background fractions, the efficiency map and background modeling. The amplitude model uncertainties arise from limited knowledge of the fixed parameters of the resonance mass models, and possible spin-2 components neglected in the baseline fit. Systematic uncertainties due to experimental effects are found to be subdominant. The breakdown of the systematic uncertainties of observables is detailed in Table 1 of the End Matter. Results of the fitted parameters with their uncertainties are also summarized in Table 2 of the End Matter.
@@
The systematic uncertainty due to fixed parameters in the amplitude model is estimated using repeated fits to the data. Each time the masses, widths, and hadron radii of intermediate resonances are varied within their uncertainties and an amplitude fit is performed under this model. The standard deviation of the distribution of the fit results among the repeated fits is taken as the corresponding systematic uncertainty. For the $B\rightarrow VV$ decay, the relative orbital angular momentum of the two vector mesons, $L$, could take the values $L=0$ or $L=2$ for the $A_{0}$ and $A_{\parallel}$ amplitudes, and $L=1$ for the $A_{\perp}$ amplitude. Various $L$ configurations are studied and the results are compared with that of the baseline fit with $L=0$, and the largest difference is taken as the systematic uncertainty. The scalar component of the ${{\pi}^{+}}{{\pi}^{-}}$ mass spectrum can also be modeled using the K-matrix approach [74], as detailed in [75]. The difference between the results obtained with this alternative model and those from the baseline fit using explicit $f_{0}$ resonances is taken as a systematic uncertainty. Furthermore, possible contributions of spin-2 resonances to the decay are assessed by including the $f_{2}(1270)\rightarrow{{\pi}^{+}}{{\pi}^{-}}$ or $K_{2}^{*}(1430)^{+}\rightarrow{{K}^{0}_{\mathrm{S}}}{{\pi}^{+}}$ components in the fit. The systematic uncertainties are estimated using the mean value of deviations, between the fits using extended model and baseline model, over an ensemble of pseudoexperiments.
@@
$f_{L}^{+}\equiv\frac{|A_{0}|^{2}}{\sum_{\lambda}|A_{\lambda}|^{2}}={0.491\pm 0.083\pm 0.025},\quad f_{L}^{-}\equiv\frac{|\kern 1.79993pt\overline{\kern-1.79993ptA}_{0}|^{2}}{\sum_{\lambda}|\kern 1.79993pt\overline{\kern-1.79993ptA}_{\lambda}|^{2}}={0.794\pm 0.025\pm 0.006}.$
@@
$\mathcal{A}_{{C\!P}}\equiv\frac{\sum_{\lambda}(|\kern 1.79993pt\overline{\kern-1.79993ptA}_{\lambda}|^{2}-|A_{\lambda}|^{2})}{\sum_{\lambda}(|\kern 1.79993pt\overline{\kern-1.79993ptA}_{\lambda}|^{2}+|A_{\lambda}|^{2})}={0.507\pm 0.062\pm 0.024}.$
@@
$\mathcal{A}_{{C\!P}}(A_{0})$ $\equiv\frac{|\kern 1.79993pt\overline{\kern-1.79993ptA}_{0}|^{2}-|A_{0}|^{2}}{|\kern 1.79993pt\overline{\kern-1.79993ptA}_{0}|^{2}+|A_{0}|^{2}}$ $={0.664\pm 0.083\pm 0.029},$
@@
$\mathcal{A}_{{C\!P}}(A_{\parallel})$ $={-0.063\pm 0.137\pm 0.027},\phantom$$\mathrm${rad}}}
@@
The $\it{true}$ TPA $\mathcal{A}_{\it{true}}^{(1)}$ deviates from zero with a significance of approximately four standard deviations, indicating evidence of CPC\!P violation in interference. The $\it{fake}$ TPA $\mathcal{A}_{\it{fake}}^{(1)}$ is nonzero with a significance of about six standard deviations, marking parity violation in interference [78].
In summary, an amplitude analysis of the decay ${{{B}^{+}}}\rightarrow({{\pi}^{+}}{{\pi}^{-}})({{K}^{0}_{\mathrm{S}}}{{\pi}^{+}})$ is performed within the mass ranges $0.30<m_{{{\pi}^{+}}{{\pi}^{-}}}<1.10\text{\,Ge\kern-1.00006ptV\!/}c^{2}$ and $0.75<m_{{{K}^{0}_{\mathrm{S}}}{{\pi}^{+}}}<1.20\text{\,Ge\kern-1.00006ptV\!/}c^{2}$. The longitudinal polarization and CPC\!P asymmetries are studied for the ${{{B}^{+}}}\rightarrow\rho(770)^{0}{{K}^{*}}(892)^{+}$ decay. The CPC\!P-averaged longitudinal fraction is measured to be $f_{L}={0.720\pm 0.028\text{\,(stat)}\pm 0.009\text{\,(syst)}}$. The direct CPC\!P asymmetry is determined to be $\mathcal{A}_{{C\!P}}={0.507\pm 0.062\text{\,(stat)}\pm 0.024\text{\,(syst)}}$, and the significance of CPC\!P violation is measured to exceed nine standard deviations using a likelihood-ratio test. This is the first observation of CPC\!P violation in the ${{{B}^{+}}}\rightarrow\rho(770)^{0}{{K}^{*}}(892)^{+}$ decay. These measurements are consistent with those previously reported by the BaBar collaboration [37], with much better precision. The CPC\!P asymmetries are also derived for the magnitudes and phases of the three subamplitudes, suggesting that the longitudinal component drives the total CPC\!P violation. These results help to constrain theoretical models used to calculate $B$-hadron decays, shed light on the dynamics of heavy-flavor decays and contribute to the understanding of the polarization puzzle of $B$-meson decays.